# Analysis and Simulation of a Fluid-Heat System in a Thin, Rough Layer in Contact With a Solid Bulk Domain

### v2

@@ Abstract
We investigate the effective coupling between heat and fluid dynamics within a thin fluid layer in contact with a solid structure via a rough surface. Moreover, the opposing vertical surfaces of the thin layer are in relative motion. This setup is particularly relevant to grinding processes, where cooling lubricants interact with the rough surface of a rotating grinding wheel. The resulting model is nonlinearly coupled through $(i)$ temperature-dependent viscosity and $(ii)$ convective heat transport. The underlying geometry is highly heterogeneous due to the thin rough surface characterized by a small parameter $\epsilon>0$ that represents both the height of the layer and the periodicity of the roughness.
@@
Keywords: Homogenization, mathematical modeling, dimension reduction, two-scale convergence, thin fluid films, numerical simulations
MSC 2020: 35B27, 76-10, 76M50, 65N30, 80M40, 80A19 {NoHyper} 1Corresponding author, Email: tomfre@uni-bremen.de
@@ 1 Introduction
This research is motivated by grinding processes, which typically involve three main components: the grinding wheel, the workpiece, and a cooling lubricant. Here, the workpiece is the material that is processed or shaped through grinding and the cooling lubricant is a fluid that is used to reduce friction and remove heat during the grinding process [38]. The grinding wheel itself is usually a composite material consisting of abrasive grains (e.g., diamonds) and a cementing matrix called bond. Given the complex granular surface of the grinding wheel, thin gaps and channels emerge in the contact area of the wheel and the workpiece, allowing the coolant to flow through. Understanding the formation and the influence of these microstructures is crucial in optimizing the usage of coolants [54, 56]. Naturally, resolving the granular wheel structure and the potentially rough workpiece surface is numerically not feasible, creating a need for homogenized models that account for microscopic effects in an effective way.
Macroscopic models describing the heat dynamics in the grinding gap via an interface temperature field can be found in the literature; see [33] and the references therein. These models are heuristically derived and the microstructures and fluid flow in the gap are not accounted for. Additionally, fluid pressure estimates are sometimes also investigated and used as a tool to control the supply of the coolant fluid [45, 58]. The fluid dynamics in the gap is strongly influenced by the viscosity of the coolant, which in reality is temperature dependent [30, 50]; and given the large temperature variations during the grinding process, this effect should not be ignored as it can influence the quality of the results [18, 57]. Numerical simulations to predict the fluid pressure in the contact zone are usually done with the help of the Reynolds equation [53, 55], but often without taking into account the temperature variations.
@@
Closely related to grinding are lubrication problems. Here, the goal is not to remove material from an object, but instead to reduce the friction between two moving objects with the help of oils. Often, temperature variations are encountered (either by friction-induced heat or external supply) that influence the viscosity of the fluid [9, 35]. Another related area is fractured media, where thin cracks are present in a larger bulk domain. Solutes and/or heat can be transported through these thin cracks, see [40] and the references therein. We also point to [39] for a numerical study on temperature-dependent viscosity on flows through fractures. Although there are differences, one typically expects much lower velocities than in our case, the analytical results presented for the temperature could be transferred.
@@
Starting with an idealized periodic setup with a small parameter $\varepsilon\ll 1$ characterizing both the length scale of the periodicity as well as the average thickness of the grinding gap, we rigorously derive an effective, homogenized model accounting for the nonlinear coupling between heat and fluid flow dynamics in such a setup. This is accomplished by conducting a limiting procedure $\varepsilon\to 0$ within the framework of two-scale homogenization for thin domains.
@@
From a mathematical point of view, the problem consists of two main aspects: a large bulk domain where the rough interface acts similar to a rough boundary and a thin rough layer. For bulk domains with a rough boundary, homogenization procedures, including diffusion or fluid flow problems, can be found in [10, 16, 44], or in the case of rough interfaces we point to [19]. The case where the amplitude of the roughness does not scale with $\varepsilon$ has also been considered in the literature, see for example [16, 19, 47] and the references therein. Additionally, there are some connections to dimension reductions in fractured media, where inside a bulk domain a thin layer of width $\varepsilon$ with different material properties appears (for example, a locally varying permeability in a porous medium). Both formal as well as rigorous limit procedures have been carried out for such problems and the thin layer reduces to effective interface equations, see [2, 40, 49].
@@
On the other hand, homogenization of diffusion problems in thin layers with periodic structures has been extensively studied before with the concept of two-scale convergence; see for example [29, 46]. Effective fluid flow through thin layers was rigorously investigated in [11] and later extended for a periodic roughness in [12]. Starting with the stationary Stokes equation, they proved the validity of the Reynolds equation [51] when the layer height is small when compared to the length of the roughness period. If the periodicity and the height are of the same order (usually called the Stokes roughness regime), a model similar to a Darcy equation is derived. The Reynolds equation plays an important role in lubrication, which has a setup similar to our underlying scenario. The equation has been further studied and homogenized in the literature [6, 13, 17, 23]; see also [42, Section 1] for a good overview. For layers without roughness, the first derivation of the Reynolds equation with the concept of two-scale convergence was achieved in [43]. Flow through a thin layer, without oscillations, coupled with a temperature-dependent viscosity was studied in [5], where a Reynolds equation with temperature-dependent viscosity was derived via asymptotic expansion.
@@
To be more specific, we consider a complex geometry incorporating a rough, thin layer similar to the setup in [12], where the boundaries of the layer are additionally in relative motion to each other (see Fig. 1). In this geometry, we are considering the coupling between a quasi-stationary Stokes system for the fluid flow and the heat dynamics. Here, strong coupling comes into play via the convection and temperature-dependent viscosity. We assume that the viscosity is a Lipschitz continuous function with respect to the temperature. We show that the resulting nonlinear problem has a weak solution (Theorem 1) via a Schauder fixed point argument where uniqueness follows under additional regularity assumptions at the solution(Theorem 2). In particular, any strong solution is necessarily unique. We then perform a limiting procedure $\varepsilon\to 0$ in the framework of two-scale convergence for thin domains and, as our main result, arrive at a homogenized model which captures the effective interaction between heat dynamics and fluid flow (Theorem 3). A novel aspect of our work is the derivation of $\varepsilon$-independent $L^{\infty}$ bounds of the temperature field, utilizing results of [41]. These are used to obtain strong two-scale convergence in the thin layer, which is required to pass to the limit in the nonlinear terms. To this end, we also show in Section 3.1 multiple ($\varepsilon$-dependent) trace estimates and embedding estimates for both the thin layer and the bulk domain. Please note that the fluid part of this effective model (e.g., when the temperature is constant) is the same as the limit presented in [12, Section 3.3]. Similarly, as for the $\varepsilon$-problem, we are able to establish uniqueness for the homogenized model under additional regularity assumptions (Theorem 4).
@@
A small convergence study $\varepsilon\to 0$ is carried out numerically in two dimensions, investigating the order of convergence of the error between the effective and microscale model averaged over the thin layer. Moreover, we conduct several numerical experiments showcasing and investigating the effects of the temperature-dependent viscosity, the roughness pattern, as well as the inflow condition. The code of our implementation is uploaded and made freely accessible on GitHub [24].
@@
2 Underlying geometry, model and assumptions
2.1 Geometry setup
@@
The general setup considered is presented in Fig. 1. We consider spatial domains of dimension $d=2,3$. The unit vectors are denoted with $e_{i}\in\mathbb{R}^{d}$ for $i=1,\dots,d$ and the $i$-dimensional unit cube will be denoted by $Y^{i}=(0,1)^{i}$. The overall domain $\Omega\subset\mathbb{R}^{d}$ is assumed to be of the form $\Omega=\Sigma\times(0,H)$ where $\Sigma\subset\mathbb{R}^{d-1}$ is a bounded $d-1$ dimensional cuboid $\Sigma\subset\mathbb{R}^{d-1}$ with corner coordinates in $\mathbb{Z}^{d-1}$. [^2] So $\Omega$ is also a $d$-dimensional cuboid. By a slight abuse of notation, we will understand $\Sigma$ both as a set in $\mathbb{R}^{d-1}$ and the lower boundary part $\Sigma\times\{0\}\subset\partial\Omega$. Any point $x\in\Omega$ will also be denoted by $x=(\tilde{x},x_{d})\subset\Sigma\times(0,H)$, the same notation is used for $y\in Y^{d}$.
@@
The rough interface between solid and fluid is given through a $Y^{d-1}$-periodic $C^{1}$ function $\gamma:\overline{Y^{d-1}}\to[\gamma_{0},1]$ for some $10<\gamma_{0}<$ 1. This lower bound $\gamma_{0}>0$ ensures that the interface does not touch the bottom boundary; see Remark 2 for a discussion regarding this choice and case $\gamma_{0}=0$ and a connection to our motivating application. We introduce the reference cell $Z\subseteq Y^{d}$ and the reference interface $\Gamma$ given by
$Z=\left\{x\in Y^{d}:x_{d}<\gamma(\tilde{x})\right\}\quad\text{ and }\quad\Gamma=\left\{x\in Y^{d}:x_{d}=\gamma(\tilde{x})\right\}.$
Since the corners of $\Sigma$ are at integer coordinates, it can be tiled perfectly with $\varepsilon Y^{d-1}$ for $\varepsilon:=\varepsilon_{n}=n^{-1}$ for any $n\in\mathbb{N}$. In the following, we suppress the $n$ and note that any limit $\varepsilon\to 0$ is understood as a limit $n\to\infty$. With that in mind, we introduce the rough solid domain $\Omega^{s}_{\varepsilon}$ and the rough, thin fluid layer $\Omega^{f}_{\varepsilon}$ by (assuming that $\varepsilon<H$)
$\Omega^{s}_{\varepsilon}=\left\{x\in\Omega:x_{d}>\varepsilon\gamma\left(\nicefrac{{\tilde{x}}}{{\varepsilon}}\right)\right\}\quad\text{and}\quad\Omega^{f}_{\varepsilon}=\left\{x\in\Omega:x_{d}<\varepsilon\gamma\left(\nicefrac{{\tilde{x}}}{{\varepsilon}}\right)\right\}.$
Here, roughness refers to the fast, oscillatory structure (due to $\nicefrac{{x}}{{\varepsilon}}$) of the interface between the fluid and grinding wheel which is given by
$\Gamma_{\varepsilon}=\left\{x\in\Omega:x_{d}=\varepsilon\gamma\left(\nicefrac{{\tilde{x}}}{{\varepsilon}}\right)\right\}.$
Note that $\Omega=\Omega^{s}_{\varepsilon}\cup\Omega^{f}_{\varepsilon}\cup\Gamma_{\varepsilon}$. We denote the outer boundary of the solid domain by $\Sigma^{s}_{\varepsilon}=\partial\Omega^{s}_{\varepsilon}\setminus\Gamma_{\varepsilon}$. Now, let $\Sigma^{\text{in}},\Sigma^{\text{out}}\subset\partial\Sigma$ be two disjoint sets with positive surface measure such that $\Sigma^{\text{in}}\cup\Sigma^{\text{out}}=\partial\Sigma$. We then split the boundary of $\Omega^{f}_{\varepsilon}$ into
@@
$\Sigma^{\text{in}}_{\varepsilon}=\left\{(\tilde{x},x_{d}):\tilde{x}\in\Sigma^{\text{in}},x_{d}\in\left(0,\varepsilon\gamma\left(\nicefrac{{\tilde{x}}}{{\varepsilon}}\right)\right)\right\}\quad\text{ and }\quad\Sigma^{\text{out}}_{\varepsilon}=\left\{(\tilde{x},x_{d}):\tilde{x}\in\Sigma^{\text{out}},x_{d}\in\left(0,\varepsilon\gamma\left(\nicefrac{{\tilde{x}}}{{\varepsilon}}\right)\right)\right\}.$
@@
In our motivating example of a grinding process, the grinding wheel moves over the workpiece, with a given velocity $u_{\text{motion}}$. Since we consider only the section of the grinding gap, it is reasonable to assume that $u_{\text{motion}}\in\mathbb{R}^{d}$ and $u_{\text{motion}}\cdot e_{d}=0$. Because of the movement both domains $\Omega^{s}_{\varepsilon}$ and $\Omega^{f}_{\varepsilon}$ would be time-dependent. To circumvent the homogenization of a moving domain we will view the problem from the reference frame of the grinding wheel and obtain a fixed geometry, see also the model description in the next section.
Remark 1.
In the case of grinding, the lower boundary of the fluid layer would also be given by a rough surface describing the workpiece. In our setup, this roughness may be characterized by another function $\gamma_{2}$, but with the same period of $\varepsilon$. It is also reasonable to assume that the workpiece roughness is constant in the direction of $u_{\text{motion}}$, therefore one can again work in a fixed reference frame. To simplify the notation we only consider one oscillating interface, but the following analysis can be transferred to the more general case. The smoothness assumption of $\gamma$ stems from technical reasons and is needed to estimate the fluid pressure.
With the subscript ##\## we will indicate a space of periodic functions in the directions $11,\dots,d-$ 1, for example
$H^{1}_{\#}(Z)\coloneqq\left\{u\in H^{1}_{\text{loc}}(\mathbb{R}^{d}):u_{|Z}\in H^{1}(Z),u(y+e_{i})=u(y)\text{ for almost all }y\in Z\text{ and }i=1,\dots,d-1\right\}.$
Additionally, we define $\nabla_{\tilde{x}}u=(\partial_{x_{1}}u,\dots,\partial_{x_{d-1}}u,0)$, where a 00 is appended such that the gradient it compatible with the usual gradient in $\mathbb{R}^{d}$. With $\|\cdot\|_{H^{1}(\Omega^{f}_{\varepsilon}\cup\Omega^{s}_{\varepsilon})}$ we will denote the combined $H^{1}$ norm on both subdomains, e.g.,
@@
2.2 Underlying model equations
For a function $\varphi\in L^{2}(\Omega)$, we use the notation $\varphi_{\varepsilon}^{k}:=\varphi_{|\Omega_{\varepsilon}^{k}}$ ($k=f,s$) for the restriction of $\varphi$ to the subdomain $\Omega_{\varepsilon}^{k}$. A similar notation holds for the coefficients of the models, which are assumed to be constant in each subdomain. We suppress superscripts and the index $\varepsilon$ when possible to improve readability.
We consider a reference frame of the grinding wheel, where instead of a moving interface $\Gamma_{\varepsilon}$ we have a prescribed velocity at the bottom boundary $\Sigma$. To describe the temperature $\theta_{\varepsilon}=(\theta^{s}_{\varepsilon},\theta^{f}_{\varepsilon})$, we utilize heat equations in both domains
[…]
Here, $\kappa=k_{f}$ in $\Omega^{f}_{\varepsilon}$ and $\kappa=\kappa_{s}$ in $\Omega^{s}_{\varepsilon}$ are the heat conductivities in each subdomain and $u_{\varepsilon}$ is the fluid velocity given via a Stokes system below. As they have no specific bearing on the mathematical analysis or the homogenization result, we have set the volumetric heat capacity to 1 in both phases.

### v3

@@ Abstract
We investigate the effective coupling between heat and fluid dynamics within a thin fluid layer in contact with a solid structure via a rough surface. Moreover, the opposing vertical surfaces of the thin layer are in relative motion. This setup is motivated by grinding processes, where cooling lubricants interact with the rough surface of a rotating grinding wheel. The resulting model is nonlinearly coupled through $(i)$ temperature-dependent viscosity and $(ii)$ convective heat transport. The underlying geometry is highly heterogeneous due to the thin rough surface characterized by a small parameter $\epsilon>0$ that represents both the height of the layer and the periodicity of the roughness.
@@
keywords:
Homogenization, mathematical modeling, dimension reduction, two-scale convergence, thin fluid films, numerical simulations
@@ 1 Introduction
This research is motivated by grinding processes, which typically involve three main components: the grinding wheel, the workpiece, and a cooling lubricant. Here, the workpiece is the material that is processed or shaped through grinding and the cooling lubricant is a fluid that is used to reduce friction and remove heat during the grinding process [1]. Given the complex granular surface of the grinding wheel, thin gaps and channels emerge in the contact area of the wheel and the workpiece, allowing the coolant to flow through. Understanding the formation and the influence of these microstructures is crucial in optimizing the usage of coolants [2, 3]. Naturally, resolving the granular wheel structure and the potentially rough workpiece surface is numerically not feasible, creating a need for homogenized models that account for microscopic effects in an effective way. Macroscopic models describing the heat dynamics in the grinding gap via an interface temperature field can be found in the literature; see [4] and the references therein. These models are heuristically derived and the microstructures and fluid flow in the gap are not accounted for. The fluid dynamics in the gap is strongly influenced by the viscosity of the coolant, which in reality is temperature dependent [5, 6]; and given the large temperature variations during the grinding process, this effect should not be ignored as it can influence the quality of the results [7, 8]. Numerical simulations to predict the fluid pressure in the contact zone are usually done with the help of the Reynolds equation [9, 10], but often without taking into account the temperature variations.
@@
Similar setups appear in lubrication problems and bearings. There, the aim is not to remove material, but instead to reduce the friction between two moving objects with the help of oils. Often, temperature variations are encountered (either by friction-induced heat or external supply) that influence the viscosity of the fluid [11, 12]. Another related area is fractured media, where thin cracks are present in a larger bulk domain. Solutes and/or heat can be transported through these thin cracks, see [13] and the references therein. We also point to [14] for a numerical study on temperature-dependent viscosity on flows through fractures. Although there are differences, one typically expects much lower velocities than in our case, the analytical results presented in this article for the temperature could still be transferred.
We point out that while our model is motivated by a grinding process, we do not try to capture all or even most of the relevant aspects of such a process; for example, we do not include solid-to-solid contact, time-dependent contact zones, or mechanical deformations. Instead, we consider a simplified framework to study some phenomena in a mathematically rigorous setting. We acknowledge that our assumptions are strong idealizations and may not directly reflect a real-world application. Many of the not included aspects are more challenging to model mathematically, and remain open from a rigorous mathematical standpoint. However, we believe that the mathematical results in this article still offer a starting point for further studies and are also relevant to other problems.
@@
Starting with an idealized periodic setup with a small parameter $\varepsilon\ll 1$ characterizing both the length scale of the periodicity as well as the average thickness of the thin layer, we rigorously derive an effective, homogenized model accounting for the nonlinear coupling between heat and fluid flow dynamics. This is accomplished by conducting a limiting procedure $\varepsilon\to 0$ within the framework of two-scale homogenization for thin domains.
@@
From a mathematical point of view, the problem consists of two main aspects: a large bulk domain where the rough interface acts similarly to a rough boundary, and a thin rough layer. For bulk domains with a rough boundary, homogenization procedures, including diffusion or fluid flow problems, can be found in [15, 16, 17], or in the case of rough interfaces, we point to [18]. The case where the amplitude of the roughness does not scale with $\varepsilon$ has also been considered in the literature, see for example [16, 18, 19] and the references therein. Additionally, there are some connections to dimension reductions in fractured media, where inside a bulk domain a thin layer of width $\varepsilon$ with different material properties appears (for example, a locally varying permeability in a porous medium). Both formal as well as rigorous limit procedures have been carried out for such problems and the thin layer reduces to effective interface equations, see [20, 13, 21].
@@
On the other hand, homogenization of diffusion problems in thin layers with periodic structures has been extensively studied before with the concept of two-scale convergence; see [22, 23]. Effective fluid flow through thin layers was rigorously investigated in [24] and later extended for a periodic roughness in [25]. Starting with the stationary Stokes equation, they proved the validity of the Reynolds equation [26] when the layer height is small compared to the length of the roughness period. If the periodicity and the height are of the same order (usually called the Stokes roughness regime), a model similar to a Darcy equation is derived. The Reynolds equation plays an important role in lubrication, which has a setup similar to our underlying scenario. The equation has been further studied and homogenized in the literature [27, 28, 29, 30]; see also [31, Section 1] for a good overview. For layers without roughness, the first derivation of the Reynolds equation with the concept of two-scale convergence was achieved in [32]. Flow through a thin layer, without oscillations, coupled with a temperature-dependent viscosity was studied in [33], where a Reynolds equation with temperature-dependent viscosity was derived via asymptotic expansion. Furthermore, fluid flow in thin (porous) layers has been rigorously homogenized with the concept of two-scale convergence [34, 35] and flow through a porous interface [36].
@@
To be more specific, we consider a geometry incorporating a rough, thin layer, where the boundaries of the layer are additionally in relative motion to each other. In this geometry, we are considering the coupling between a quasi-stationary Stokes system for the fluid flow and the heat dynamics. Here, strong coupling comes into play via the convection and temperature-dependent viscosity. We assume that the viscosity is a Lipschitz continuous function with respect to the temperature. We show that the resulting nonlinear problem has a weak solution (Theorem 1) via a Schauder fixed point argument. A priori estimates of the temperature are obtained with energy estimates. The pressure is estimated with the help of Bogovskii operators for thin layers, which can be found in existing literature [34, 35]. Uniqueness follows under additional regularity assumptions at the solution (Theorem 2). In particular, any strong solution is necessarily unique.
The main novel aspect of our work is the derivation of $\varepsilon$-independent $L^{\infty}$ bounds of the temperature field in the thin layer (Lemma 6), utilizing results of [37]. These are used to obtain strong two-scale convergence, which is required to pass to the limit in the nonlinear terms. To this end, we also show in Section 3.1 multiple ($\varepsilon$-dependent) trace estimates and embedding estimates for both the thin layer and the bulk domain. The limit procedure itself can then be carried out with arguments already established in prior works [38, 36, 35], with slight modifications of the arguments to obtain a two-pressure representation and the boundary conditions. The homogenized model is stated in Theorem 4. Similarly, as for the $\varepsilon$-problem, we are able to establish uniqueness for the homogenized model under additional regularity assumptions (Theorem 5).
@@
In the end, a small convergence study $\varepsilon\to 0$ is carried out numerically, investigating the order of convergence of the error between the effective model and the microscale model. Moreover, we conduct several numerical experiments showcasing and investigating the effects of the temperature-dependent viscosity, the roughness pattern, as well as the inflow condition. The code of our implementation is uploaded and made freely accessible on GitHub [39].
The limit model (see Theorem 4) consists of a heat equation for the solid bulk domain coupled to heat and fluid dynamics on a lower-dimensional surface. The effective equation for the fluid temperature recovers the models derived for diffusion problems in thin connected layers [22, 36], but with an additional advection term induced by the effective flow field. The effective fluid velocity is given by a two-pressure Stokes system, which can be reduced to a Darcy-like equation with a temperature-dependent viscosity. Ignoring the viscosity, the same effective equation as in [25] is obtained; with fewer mathematical assumptions thanks to the use of two-scale convergence. The effective flow model is also similar to [35] where pressure-driven flows in thin porous layers were analyzed. The cell problems (4.2.2) used to compute the effective permeability in the layer are the same as in classical porous media [40, Eq. (1.11)] and also appear, with slightly modified boundary conditions, in [35, Eq. (6.4)]. In addition to the permeability, a term induced by the relative motion appears in the Darcy equation, see Section 4.2.2. A similar cell problem occurs in [36, Eq. (20)] when computing the effective flow through a thin porous interface. We conclude that individual parts of the derived limit model have appeared in the literature in different related setups. The main novelty of this work is to combine these interesting features, in particular with respect to the coupling through the temperature-dependent viscosity and the advection term. Mathematically speaking, this introduces several difficulties. For one, the nonlinear coupling necessitates a careful fixed-point argument to ensure existence of solutions where uniqueness is generally only to be expected under higher regularity assumptions (see Theorem 2). In addition, uniform (w.r.t. the scale parameter $\varepsilon$) $L^{\infty}$-estimates are needed for this fixed-point argument to work. Due to the domain interactions (bulk domain connected to a thin fluid layer), this calls for the careful use of $\varepsilon$-controlled embeddings (see lemmas 3 and 6).
From an application point of view, this is of particular interest whenever large temperature variations are possible as this generally implies changes in viscosity. Returning specifically to the application of grinding, a similar model was heuristically introduced in [4]. There, the coolant is also modeled by a function defined on the interface between the wheel and the workpiece. We could therefore give some mathematical justification for the model used in the literature. However, the heuristic model is not completely the same; the fluid velocity is predetermined, no effective parameters depending on the microstructure appear, and the heat diffusion is set to zero along the interface.
@@
2 Problem formulation and assumptions
In this section, we provided the setup of the underlying geometry and the concrete differential equations on the micro scale.
2.1 Description of the geometry
@@
The general setup is presented in Fig. 1. We consider spatial domains of dimension $d=2,3$. The unit vectors are denoted with $e_{i}\in\mathbb{R}^{d}$ for $i=1,\dots,d$ and the $i$-dimensional unit cube is denoted by $Y^{i}=(0,1)^{i}$. The overall domain $\Omega\subset\mathbb{R}^{d}$ is assumed to be of the form $\Omega=\Sigma\times(0,H)$ where $\Sigma\subset\mathbb{R}^{d-1}$ is a bounded d1d-1 dimensional cuboid $\Sigma\subset\mathbb{R}^{d-1}$ with corner coordinates in $\mathbb{Z}^{d-1}$. [^1] So $\Omega$ is also a $d$-dimensional cuboid. By a slight abuse of notation, we understand $\Sigma$ both as a set in $\mathbb{R}^{d-1}$ and the lower boundary part $\Sigma\times\{0\}\subset\partial\Omega$. Any point $x\in\Omega$ is also denoted by $x=(\tilde{x},x_{d})\in\Sigma\times(0,H)$, the same notation is used for $y\in Y^{d}$.
@@
To construct the rough thin layer, we introduce the open periodic reference cell $Z\subseteq Y^{d}$ for the fluid domain. We assume, for technical reasons, that
  • $Z$ is Lipschitz and connected, $Y^{d}\setminus Z$ is either empty or also Lipschitz and connected,
  • there is a $\gamma_{0}>0$ such that $y_{d}\geq\gamma_{0}$ for all $y\in Y^{d}\setminus Z$,
  • the domain is periodic in the lateral directions, meaning
    $\partial Z\cap\{y_{i}=0\}+e_{i}=\partial Z\cap\{y_{i}=1\},\quad\text{for }i=1,\dots,d-1,$
  • if $Y^{d}\setminus Z\neq\emptyset$, then $Y^{d}\setminus Z$ touches the top boundary:
    $\left|\partial(Y^{d}\setminus Z)\cap\{y_{d}=1\}\right|>0.$
The lower bound $\gamma_{0}>0$ ensures that the solid does not touch the bottom boundary. The reference interface $\Gamma$ and the combined upper and lower interface of $Z$, denoted by $\Gamma_{d}$, are given by
$\Gamma\coloneqq\operatorname{int}(\partial Z\cap\partial(Y\setminus Z)),\quad\Gamma_{d}:=\Gamma\cup\{y_{d}=0\}.$
Since the corners of $\Sigma$ are at integer coordinates, we can tile $\Sigma$ with $\varepsilon Y^{d-1}$ for $\varepsilon\coloneqq\varepsilon_{n}=n^{-1}$ for any $n\in\mathbb{N}$. In the following, we suppress the $n$ and note that any limit $\varepsilon\to 0$ is understood as a limit $n\to\infty$. With that in mind, we define $\mathcal{K}_{\varepsilon}\coloneqq\{k\in\mathbb{Z}^{d-1}\times\{0\}:\varepsilon(Y+k)\subset\Omega\}$. We can introduce the rough, thin fluid layer $\Omega^{f}_{\varepsilon}$, the rough solid domain $\Omega^{s}_{\varepsilon}$, and the rough interface $\Gamma_{\varepsilon}$ by (assuming that $\varepsilon<H$)
$\Omega^{f}_{\varepsilon}\coloneqq\operatorname{int}\left(\bigcup_{k\in\mathcal{K}_{\varepsilon}}\varepsilon(\overline{Z}+k)\right),\quad\Omega^{s}_{\varepsilon}\coloneqq\Omega\setminus\overline{\Omega}_{\varepsilon}^{f},\quad\text{and}\quad\Gamma_{\varepsilon}\coloneqq\operatorname{int}\left(\partial\Omega^{f}_{\varepsilon}\cap\partial\Omega^{s}_{\varepsilon}\right).$
By construction, $\Omega^{f,s}_{\varepsilon}$ are connected, and we assume they are both also Lipschitz. Note that $\Omega=\Omega^{s}_{\varepsilon}\cup\Omega^{f}_{\varepsilon}\cup\Gamma_{\varepsilon}$. We denote the outer boundary of the solid domain by $\Sigma^{s}_{\varepsilon}=\partial\Omega^{s}_{\varepsilon}\setminus\Gamma_{\varepsilon}$. Now, let $\Sigma^{\text{in}},\Sigma^{\text{out}}\subset\partial\Sigma$ be two disjoint sets with positive surface measure such that $\Sigma^{\text{in}}\cup\Sigma^{\text{out}}=\partial\Sigma$. We then split the boundary of $\Omega^{f}_{\varepsilon}$ into
@@
$\Sigma^{\text{in}}_{\varepsilon}\coloneqq(\Sigma^{\text{in}}\times(0,\varepsilon))\cap\partial\Omega^{f}_{\varepsilon}\quad\text{ and }\quad\Sigma^{\text{out}}_{\varepsilon}\coloneqq(\Sigma^{\text{out}}\times(0,\varepsilon))\cap\partial\Omega^{f}_{\varepsilon}.$
@@
With $n$, we denote the outer normal vector on the domains $\Omega,\Omega^{f,s}_{\varepsilon}$ and $Z$. On the common interface $\Gamma_{\varepsilon}$ we further write $n^{f,s}$ to denote the outer normal vector corresponding to $\Omega^{f,s}_{\varepsilon}$, but the superscript is suppressed whenever the direction is clear from context. Additionally, the normal vector $n_{\Sigma}$ is introduced in Equation 2 for the lateral boundaries of the fluid domain.
With the subscript #\#, we indicate a space of periodic functions in the directions $1,\dots,d-1$; in particular, we use
$C_{\#}(Z)$ $\coloneqq\left\{u\in C(\mathbb{R}^{d})\ :u(x+e_{i})=u(x)\text{ for all }x\in\mathbb{R}^{d}\text{ and }i=1,\dots,d-1\right\},$
$L^{2}_{0,\#}(Z)$ $\coloneqq\left\{u\in L^{2}_{\text{loc}}(\mathbb{R}^{d})\ :\ u_{|Z}\in L^{2}_{0}(Z),\ u(x+e_{i})=u(x)\text{ for a.a. }x\in\mathbb{R}^{d}\text{ and }i=1,\dots,d-1\right\},$
$H^{1}_{\#}(Z)$ $\coloneqq\left\{u\in H^{1}_{\text{loc}}(\mathbb{R}^{d}):u_{|Z}\in H^{1}(Z),u(x+e_{i})=u(x)\text{ for a.a. }x\in\mathbb{R}^{d}\text{ and }i=1,\dots,d-1\right\},$
$H^{1}_{\#,\Gamma_{d}}(Z)$ $\coloneqq\left\{u\in H^{1}_{\#}(Z)\ :\ u=0\text{ for a.a. }y\in\Gamma_{d}\right\}.$
Additionally, we define $\nabla_{\tilde{x}}u=(\partial_{x_{1}}u,\dots,\partial_{x_{d-1}}u,0)$, where a 0 is appended such that the gradient it compatible with the usual gradient in $\mathbb{R}^{d}$. With $\|\cdot\|_{H^{1}(\Omega^{f}_{\varepsilon}\cup\Omega^{s}_{\varepsilon})}$ we denote the combined $H^{1}$ norm on both subdomains, e.g.,
@@
2.2 The micro model
For a function $\varphi\in L^{2}(\Omega)$, we use the notation $\varphi_{\varepsilon}^{k}\coloneqq\varphi_{|\Omega_{\varepsilon}^{k}}$ ($k=f,s$) for the restriction of $\varphi$ to the subdomain $\Omega_{\varepsilon}^{k}$. A similar notation holds for the coefficients of the models, which are assumed to be constant in each subdomain. We suppress superscripts and the index $\varepsilon$ when possible to improve readability.
To describe the temperature $\theta_{\varepsilon}=(\theta^{s}_{\varepsilon},\theta^{f}_{\varepsilon})$, we utilize heat equations in both domains
$\partial_{t}\theta_{\varepsilon}-\operatorname{div}\left(\kappa\nabla\theta_{\varepsilon}\right)$ $=0\quad\quad$ $\text{ in }S\times\Omega^{s}_{\varepsilon},$ (1a)
$\frac{1}{\varepsilon}\partial_{t}\theta_{\varepsilon}-\frac{1}{\varepsilon}\operatorname{div}\left(\kappa\nabla\theta_{\varepsilon}-u_{\varepsilon}\theta_{\varepsilon}\right)$ $=0\quad\quad$ $\text{ in }S\times\Omega^{f}_{\varepsilon},$ (1b)
$\theta_{\varepsilon}(0,\cdot)$ $=\theta_{\varepsilon,0}$ $\text{ in }\Omega.$ (1c)
Here, $\kappa=\kappa_{f}$ in $\Omega^{f}_{\varepsilon}$ and $\kappa=\kappa_{s}$ in $\Omega^{s}_{\varepsilon}$ are the heat conductivities in each subdomain and $u_{\varepsilon}$ is the fluid velocity given via a Stokes system below. As they have no specific bearing on the mathematical analysis or the homogenization result, we have set the volumetric heat capacity to 1 in both phases.